OpenStudy (anonymous):

Completely simplify any radical expressions. 5x2 = 40 a) x = 2 sqrt 3 b) x =2 sqrt 2 c) x = 2 d) x = 3

OpenStudy (anonymous):

a) x = 2 sqrt 3

OpenStudy (paxpolaris):

\[\large 5x^2=40\]\[\implies \large x^2 = \frac {40}5=8\]\[\implies \large x^2 = 8\]\[\implies \large x = \pm \sqrt8\] \[\implies \huge x = \pm 2\sqrt2\]